OpenStudy (anonymous):

which values best approximate the solutions of -1.5x^2 + 4x + 2.2 = 0? no real solution 2.2 and 3.1 <===my answer -0.5 and 3.1 -0.5 and 2.2

OpenStudy (anonymous):

I know its not 'no real solution'

OpenStudy (superdavesuper):

hmmm why u say there is no "real soluton"?

OpenStudy (superdavesuper):

b^2-4ac shows that there are two real solutions...so plz try again :)

OpenStudy (anonymous):

The option. I know there are two solutions. One of the options is 'there is no real solution'. I said that I know it's not 'no real solution'

OpenStudy (superdavesuper):

Oh ok sorry I missed read ur post! So are u familiar with solving quadratic roots by using (-b+-SQRT(b^2-4ac))/2a?

OpenStudy (anonymous):

yes

OpenStudy (superdavesuper):

so in this case a=-1.5 b=+4 c=+22 substitute them into the eqn n solve for roots

OpenStudy (whpalmer4):

are you sure you've copied the equation properly?

OpenStudy (superdavesuper):

Ooops...thanks @whpalmer4 sorry....what I had solve is actually a=-1.5 b=+4 c=+2.2 <---- c is different....

OpenStudy (anonymous):

oops nope its supposed to be 2.2

OpenStudy (whpalmer4):

okay, then your answer choice of B is not correct..

OpenStudy (anonymous):

It's C huh?

OpenStudy (whpalmer4):

you can of course check any proposed answer by substituting it into the equation. for example, x=2.2: -1.5(2.2)^2 + 4(2.2) + 2.2 = -1.5(4.84)+8.8 + 2.2 = -1.5(5) + 9 + 2 that clearly isn't going to equal 0, right? but if that was a solution, it would.

OpenStudy (whpalmer4):

yes, C is correct. -1.5(3.1)^2 + 4(3.1) + 2.2 = -1.5*9.61 + 12.4 + 2.2 = -1.5*9.61 + 14.6 which is pretty close to 0.

OpenStudy (anonymous):

Phew glad I asked! Thanks so much!

OpenStudy (whpalmer4):

I didn't check with -0.5 because that applies to two answer choices, a bunch of work for no reward :-)

OpenStudy (whpalmer4):

but for completeness, we would want to make sure it was a solution: -1.5(-0.5)^2 + 4(-0.5) + 2.2 = -(3/2)(-1/2)^2 + 4(-1/2) + 2.2 = -3/8 - 2 + 2.2 and 3/8 = 0.375 so that's -0.175 which is acceptably close for this problem. in exact form, the equation is -(3/2)x^2 + 4x + 11/5 = 0, and the solutions are x = (1/15)(20 +/- sqrt(730)) = -0.467901, 3.13457